The pressure reading in a mercury barometer is 76 cm. Convert this into N/m^2. If a barometer were made with water of density 1.0 g/cc, what would be equivalent pressure reading in cm unit?  
Answer #1

P = rho * g h = 13.6* 10^3* 9.8* 0.76

P = 1.013* 10^5 N/m^2

=====

b)

P = h * ( rho *g)

h = 1.013* 10^5 / ( 1000* 9.8)

h = 10.336 m = 1033.6 cm

=====

Comment before rate in case any doubt, will reply for sure.. goodluck